# Think-on-Graph 3.0: Efficient and Adaptive LLM Reasoning on Heterogeneous Graphs via Multi-Agent Dual-Evolving Context Retrieval

### v1

@@ 4.5 Abalation Study
Table 3: Ablation studies of MACER components and foundation model scaling. Standard ToG-3 settings incorporates all MACER components, employs the Qwen2.5-32B-instruct as the backbone LLM, and utilizes the Jina-v3-embedding model for representation encoding.
Ablation Setting HotpotQA 2WikiMultihopQA Musique Average
EM F1 EM F1 EM F1 EM F1
Standard ToG-3 settings 0.639 0.516 0.500 0.267 0.221 0.153 0.453 0.312
MACER Components Ablation
w/o Evolving Query 0.598 0.443 0.412 0.203 0.178 0.121 0.396 0.256
w/o Evolving Sub-Graph 0.613 0.472 0.458 0.234 0.203 0.138 0.425 0.281
w/o Community Node 0.641 0.519 0.487 0.259 0.216 0.148 0.448 0.309
Foundation Model Scaling Abalation
LLM Model
Qwen2.5-14B 0.573 0.469 0.453 0.231 0.198 0.134 0.408 0.278
Qwen2.5-72B 0.668 0.538 0.523 0.281 0.235 0.162 0.475 0.327
Embedding Model
Qwen3-Embed-0.6B 0.638 0.517 0.505 0.269 0.224 0.155 0.456 0.314
Qwen3-Embed-4B 0.643 0.523 0.508 0.271 0.227 0.158 0.459 0.317
Our ablation study reveals the relative importance of each MACER component for deep reasoning performance. The most significant performance degradation occurs when removing the evolving query mechanism (average performance drop of 12.6% in EM and 17.9% in F1), underscoring its critical role in complex question answering, expecially when using smaller LLMs. The iterative query decomposition enables the framework to break down multifaceted questions into tractable sub-problems, which is essential for navigating the heterogeneous graph structure. Removing subgraph refinement causes a moderate performance decrease (average drop of 6.2% in EM and 10.0% in F1), indicating its importance in adapting the knowledge structure to the specific reasoning context. Interestingly, community nodes show the smallest impact on deep reasoning tasks (a slight drop in the average EM and F1 scores), suggesting that while they contribute to performance, the chunk and triplet representations carry most of the relevant information for precise answer generation. However, in broad reasoning tasks, community nodes are essential for comprehensive coverage and diversity, highlighting the complementary roles of different node types in our heterogeneous graph architecture.
@@
The foundation model scaling analysis reveals several important patterns. First, LLM capacity has a substantially greater impact on performance than embedding model size. Scaling from Qwen2.5-14B to Qwen2.5-72B yields a 16.4% average improvement in EM scores, highlighting the critical role of reasoning capability in complex QA tasks. Second, larger embedding models provide consistent but more modest improvements. Qwen3-Embed-0.6B shows a slight average EM improvement over jina-embeddings-v3, while Qwen3-Embed-4B provides a 1.6% improvement. This suggests that while retrieval quality matters and larger embedding models contribute to better performance, the LLM’s reasoning capacity remains the primary bottleneck for complex reasoning tasks. These findings provide practical guidance for resource allocation in real-world deployments.
@@ 5 Conclusion
In this work, we introduced Think-on-Graph 3.0, a novel framework that fundamentally rethinks the paradigm of RAG for complex reasoning. By proposing the Multi-Agent Context Evolution Retrieval (MACER) mechanism and a dynamic Chunk-Triplets-Community heterogeneous graph architecture, we address critical limitations in both existing graph-based RAG methods and knowledge-graph-dependent approaches. Our comprehensive experimental evaluation demonstrates that ToG-3 achieves state-of-the-art performance across multiple challenging benchmarks. The framework’s core innovation is its dual-evolving mechanism—comprising Evolving Query and Evolving Subgraph—which dynamically refines both the query representation and the underlying graph structure throughout the reasoning process. This iterative co-evolution enables deep, multi-hop inference while preserving broad coverage across complex queries. This adaptive capability proves particularly valuable for overcoming the quality constraints of static graph construction and the domain limitations of pre-existing knowledge bases. The framework’s ability to work with light LLMs also opens possibilities for more efficient and deployable AI systems. We believe the principles established in ToG-3—dynamic graph evolution, multi-agent collaboration, and heterogeneous knowledge integration—provide a foundation for the next generation of RAG systems.
6 Limitation and Future Directions
@@
Future work will explore three promising directions for further advancement. First, we plan to extend our multi-agent evolving framework to support larger-scale and more complex knowledge-intensive tasks, such as programming assistance, financial analysis, and clinical decision-making. Second, we aim to generalize our method from text to multimodal reasoning, integrating audio, image, and video modalities to construct a world model that bridges textual knowledge with perceptual grounding. Third, inspired by human cognitive science and brain science, we will explore novel architectures that combine parametric memory with large language models to unify memory and reasoning in a seamless framework, enabling more efficient knowledge retention and tool-use capabilities akin to human intelligence.
@@
  • Appendix G: Case Study for ToG-3
  • Appendix H: Graph Visualization Examples
  • Appendix I: Theoretical Support for ToG-3
  • Appendix J: LLM Prompts
@@ Appendix A Implementation Details
We implement ToG-3 experiments with the following configuration: Data Processing: Chunk size is set to 1024 tokens with 20-token overlap between consecutive chunks to maintain contextual continuity. Multi-Agent hyperparameter: Constructor Agent extracts a maximum of 2 knowledge triplets per chunk and employs hierarchical Leiden clustering (Traag et al., 2019) with maximum cluster size of 5 for community detection. Retriever Agent retrieves top-5 most relevant nodes using hybrid vector-graph similarity matching. Reflector/Responser Agent utilizes the top-5 retrieved passages as context for answer generation. Backend Infrastructure: LLM service is based on Qwen2.5-32B-Instruct (Yang et al., 2024) deployed with vLLM (Kwon et al., 2023) engine using bfloat16 precision and prefix caching enabled and greedy-search generation method, which is more stable than the Qwen3 model in mixed reasoning mode in our task; embeddings are generated using Jina-embeddings-v3 (1024-dimensional) (Sturua et al., 2024); Our server is equipped with 8 A100 40GB cards, AMD EPYC 256-core Processor, 2TB memory, and Ubuntu 20.04.1 system. and the hybrid vector-graph storage is implemented using Neo4j community edition  [^1] https://neo4j.com/product/community-edition for efficient knowledge representation and retrieval, see Appendix.H for visualized graph example.
@@ Appendix B ToG-3 Algorithms
6:  $a_{k}\leftarrow\pi_{\text{resp}}(q,\mathcal{G}_{k},\mathcal{H}_{k})$ $\triangleright$ Response Agent generates answer
7:  $r_{k}\leftarrow\pi_{\text{ref}}^{\text{suff}}(q,\mathcal{G}_{k},a_{k})$ $\triangleright$ Reflector judges sufficiency
8:  if $r_{k}=1$ then break
9:  end if
10:  $q^{\prime}_{k}\leftarrow\pi_{\text{ref}}^{\text{evolve}}(q,\mathcal{G}_{k})$ $\triangleright$ Reflector evolves query
11:  $\mathcal{G}_{k+1}\leftarrow\pi_{\text{const}}^{{\text{evolve}}}(q^{\prime}_{k},\mathcal{G}_{k})$ $\triangleright$ Constructor evolves subgraph
12:  $\mathcal{H}_{k+1}\leftarrow\mathcal{H}_{k}\cup\{(q^{\prime}_{k},a_{k},r_{k},\mathcal{G}_{k+1})\}$
13:  $k\leftarrow k+1$
14:until $k=K$
@@ C.1 Deep Reasoning Datasets
  • HotpotQA (Yang et al., 2018): A crowdsourced question answering dataset built on English Wikipedia, comprising approximately 113K questions. Each question is constructed to require the combination of information from the introductory sections of two Wikipedia articles for answering. The dataset provides two gold paragraphs per question, along with a list of sentences identified as supporting facts necessary to answer the question. HotpotQA includes various reasoning strategies such as bridge questions (involving missing entities), intersection questions (e.g., “what satisfies both property A and property B?”), and comparison questions (comparing two entities through a common attribute). It is available in two settings: a few-shot distractor setting where models are provided with 10 paragraphs including the gold ones, and an open-domain full-wiki setting where models must retrieve relevant passages from the entire Wikipedia corpus given only the question.
  • 2WikiMultihopQA (Ho et al., 2020): A multi-hop question answering dataset that contains complex questions requiring reasoning over multiple Wikipedia paragraphs. Each question is designed to necessitate logical connections across different pieces of information to arrive at the correct answer.
  • Musique (Trivedi et al., 2022): A challenging multi-hop QA dataset containing approximately 25K 2–4 hop questions, constructed by composing single-hop questions from five existing single-hop QA datasets. It is designed to feature diverse and complex reasoning paths, requiring models to integrate information from multiple hops to generate correct answers. The dataset emphasizes comprehensive evaluation of multi-step reasoning capabilities.
@@ C.2 Broad Reasoning Datasets
The following datasets are curated from the UltraDomain (Qian et al., 2025) benchmark. The benchmark construction leverages financial reports, legal contracts, and 428 college textbooks across 18 distinct domains to evaluate model versatility and adaptability in specialized and broad application scenarios:
@@ Appendix D Baselines
  • NaiveRAG (Gao et al., 2023): A standard chunk-based retrieval baseline that segments raw texts into chunks and stores them in a vector database using text embeddings. For queries, it generates vectorized representations to directly retrieve text chunks based on semantic similarity.
  • GraphRAG (Edge et al., 2024): A graph-enhanced RAG system that utilizes an LLM to extract entities and relationships from text, representing them as nodes and edges. It generates community summaries through graph clustering and employs both local (entity-based) and global (community-based) retrieval strategies for comprehensive information access.
  • LightRAG (Guo et al., 2024): A graph-structured RAG framework that employs a dual-level retrieval system combining low-level entity retrieval with high-level knowledge discovery. It integrates graph structures with vector representations for efficient retrieval of related entities and their relationships.
  • ToG-2 (Ma et al., 2024): A knowledge graph-based framework implements a tight-coupling hybrid RAG paradigm that iteratively retrieves information from both unstructured texts and structured knowledge sources. It alternates between graph retrieval and context retrieval for in-depth knowledge exploration.
  • HippoRAG-2 (Gutiérrez et al., 2025): A non-parametric continual learning framework that leverages Personalized PageRank algorithm over an open knowledge graph constructed using LLM-extracted triples. It enhances multi-hop reasoning capabilities through sophisticated graph traversal and passage integration mechanisms.
@@ Appendix E Metrics
For Deep Reasoning Tasks, we follow standard QA evaluation practices as ToG (Sun et al., 2023; Ma et al., 2024) and HippoRAG (Gutiérrez et al., 2024; 2025):
  • Exact Match (EM): Measures the percentage of predictions that exactly match the ground truth answer.
@@
For Broad Reasoning Tasks, we adopt a multi-dimensional LLM-based evaluation approach due to the complexity and open-ended nature of these queries following LightRAG (Guo et al., 2024):
@@
The LLM-based evaluation uses GPT-4o-mini as judge, with careful attention to prompt design and answer ordering to avoid positional bias. The LLM evaluation prompt is shown in Appendix.J
@@ F.2 Result Detail in Braod Reasoning Tasks
Comprehensiveness 28.4% 71.6% 32.4% 67.6% 12.4% 87.6% 34.8% 65.2%
Diversity 19.6% 80.4% 32.0% 68.0% 9.6% 90.4% 28.4% 71.6%
Empowerment 29.4% 70.6% 32.8% 67.2% 12.4% 87.6% 38.8% 61.2%
Overall 28.7% 71.3% 33.3% 66.7% 11.2% 88.8% 36.0% 64.0%
@@
Comprehensiveness 46.8% 53.2% 49.6% 50.4% 49.6% 50.4% 51.6% 48.4%
Diversity 44.4% 55.6% 48.4% 51.6% 46.8% 53.2% 52.0% 48.0%
Empowerment 25.2% 74.8% 43.2% 56.8% 29.6% 70.4% 38.4% 61.6%
Overall 47.6% 52.4% 49.2% 50.8% 48.4% 51.6% 51.2% 48.8%
@@
Comprehensiveness 38.9% 61.1% 45.6% 54.4% 33.6% 66.4% 47.6% 52.4%
Diversity 32.0% 68.0% 42.0% 58.0% 28.0% 72.0% 39.2% 60.8%
Empowerment 40.5% 59.5% 46.0% 54.0% 33.6% 66.4% 52.0% 48.0%
Overall 39.6% 60.4% 46.0% 54.0% 32.4% 67.6% 49.6% 50.4%
@@
Comprehensiveness 24.5% 75.5% 31.6% 68.4% 21.6% 78.4% 29.6% 70.4%
Diversity 18.8% 81.2% 28.0% 72.0% 17.2% 82.8% 23.6% 76.4%
Empowerment 27.8% 72.2% 32.9% 67.1% 21.6% 78.4% 34.0% 66.0%
Overall 25.6% 74.4% 32.0% 68.0% 20.4% 79.6% 31.6% 68.4%

### v2

@@ 4.4 Abalation Study
Our ablation study reveals the relative importance of each MACER component for deep reasoning performance. The most significant performance degradation occurs when removing the evolving query mechanism (average performance drop of 12.0% in EM and 16.5% in F1), underscoring its critical role in complex question answering, expecially when using smaller LLMs. Removing subgraph refinement causes a moderate performance decrease (average drop of 6.0% in EM and 9.0% in F1), indicating its importance in adapting the knowledge structure to the specific reasoning context. Interestingly, community nodes show the smallest impact on deep reasoning tasks (a slight drop in the average EM and F1 scores), suggesting that while they contribute to performance, the chunk and triplet representations carry most of the relevant information for precise answer generation. However, in broad reasoning tasks, community nodes are essential for comprehensive coverage and diversity, highlighting the complementary roles of different node types in our heterogeneous graph architecture. Note that the reranker agent also delivers a 4.6% improvement in EM and a 10.6% improvement in F1. This is because, during multi-turn RAG processes, an excessive amount of retrieved evidence can otherwise impair the response quality of the responser agent.
@@
The foundation model scaling analysis reveals several important patterns. First, LLM capacity has a substantially greater impact on performance than embedding model size. Scaling from Qwen2.5-14B to Qwen2.5-72B yields a 15.9% average improvement in EM scores, highlighting the critical role of reasoning capability in complex QA tasks. Second, larger embedding models provide consistent but more modest improvements. Qwen3-Embed-0.6B shows a slight average EM improvement over jina-embeddings-v3, while Qwen3-Embed-4B provides a 1.7% improvement. This suggests that while retrieval quality matters and larger embedding models contribute to better performance, the LLM’s reasoning capacity remains the primary bottleneck for complex reasoning tasks. These findings provide practical guidance for resource allocation in real-world deployments.
@@ 5 Conclusion
In this work, we introduced Think-on-Graph 3.0, a novel framework that fundamentally rethinks the paradigm of RAG for complex reasoning. By proposing the Multi-Agent Context Evolution Retrieval (MACER) mechanism and a dynamic Chunk-Triplets-Community heterogeneous graph architecture, we address critical limitations in both existing graph-based RAG methods and knowledge-graph-dependent approaches. Our comprehensive experimental evaluation demonstrates that ToG-3 achieves state-of-the-art performance across multiple challenging benchmarks. This adaptive capability proves particularly valuable for overcoming the quality constraints of static graph construction and the domain limitations of pre-existing knowledge bases. The framework’s ability to work with light LLMs also opens possibilities for more efficient and deployable AI systems.
Limitations
@@
Ethical Considerations
This research focuses on improving the technical performance of knowledge-enhanced language models. This work utilizes only public benchmark datasets and adheres to strict reproducibility standards. While our framework improves text generation capability, we acknowledge potential risks of generating misleading content and note that performance may reflect biases inherent in base models. We follow the ACL ethical guidelines when conducting the research in this paper.
Information About Use Of AI Assistants
In the preparation of this work, the author used AI-assisted technology (specifically, large language models such as GPT-5 and Deepseek-V3) exclusively for text refinement purposes. The AI was employed to assist in proofreading, correcting grammatical errors, and polishing linguistic expressions to improve the clarity and readability of the manuscript.
@@
  • Appendix G: Comparison of Time and Token Consumption
  • Appendix H: Additional Baselines
  • Appendix I: Case Study for ToG-3
  • Appendix J: Graph Visualization Examples
  • Appendix K: Theoretical Support for ToG-3
  • Appendix L: LLM Prompts
@@ Appendix A Implementation Details
We implement ToG-3 experiments with the following configuration: Data Processing: Chunk size is set to 1024 tokens with 20-token overlap between consecutive chunks to maintain contextual continuity. Multi-Agent hyperparameter: Constructor Agent extracts a maximum of 2 knowledge triplets per chunk and employs hierarchical Leiden clustering (traag2019Leiden) with maximum cluster size of 5 for community detection. Retriever Agent retrieves top-5 most relevant nodes using hybrid vector-graph similarity matching. Then, the Reranker reranks the top-2 relevant evidence nodes (or triples) within this retrieved subgraph. Reflector/Responser Agent utilizes the top-2 retrieved passages as context for answer generation. Backend Infrastructure: LLM service is based on Qwen2.5-32B-Instruct (qwen2.5) deployed with vLLM (vllm) engine using bfloat16 precision and prefix caching enabled and greedy-search generation method, which is more stable than the Qwen3 model in mixed reasoning mode in our task; embeddings are generated using Jina-embeddings-v3 (1024-dimensional) (sturua2024jinaembeddingsv3); we use jina-reranker-v2 (wang2025jina-reranker-v3) as the reranker model; Our server is equipped with 8 A100 40GB cards, AMD EPYC 256-core Processor, 2TB memory, and Ubuntu 20.04.1 system. and the hybrid vector-graph storage is implemented using Neo4j community edition  [^1] https://neo4j.com/product/community-edition for efficient knowledge representation and retrieval, see Appendix.J for visualized graph example.
@@ Appendix B ToG-3 Algorithms
6:  $G_{k}\leftarrow\pi_{\text{rer}}(q,\mathcal{G}_{k-1})$ $\triangleright$ Reranker Agent rerank and select the sub-graph
7:  $a_{k}\leftarrow\pi_{\text{resp}}(q,\mathcal{G}_{k},\mathcal{H}_{k})$ $\triangleright$ Response Agent generates answer
8:  $r_{k}\leftarrow\pi_{\text{ref}}^{\text{suff}}(q,\mathcal{G}_{k},a_{k})$ $\triangleright$ Reflector judges sufficiency
9:  if $r_{k}=1$ then break
10:  end if
11:  $q^{\prime}_{k}\leftarrow\pi_{\text{ref}}^{\text{evolve}}(q,\mathcal{G}_{k})$ $\triangleright$ Reflector evolves query
12:  $\mathcal{G}_{k+1}\leftarrow\pi_{\text{const}}^{{\text{evolve}}}(q^{\prime}_{k},\mathcal{G}_{k})$ $\triangleright$ Constructor evolves subgraph
13:  $\mathcal{H}_{k+1}\leftarrow\mathcal{H}_{k}\cup\{(q^{\prime}_{k},a_{k},r_{k},\mathcal{G}_{k+1})\}$
14:  $k\leftarrow k+1$
15:until $k=K$
@@ C.1 Deep Reasoning Datasets
  • HotpotQA (yang2018hotpotqa): A crowdsourced question answering dataset built on English Wikipedia, comprising approximately 113K questions. Each question is constructed to require the combination of information from the introductory sections of two Wikipedia articles for answering. The dataset provides two gold paragraphs per question, along with a list of sentences identified as supporting facts necessary to answer the question. HotpotQA includes various reasoning strategies such as bridge questions (involving missing entities), intersection questions (e.g., “what satisfies both property A and property B?”), and comparison questions (comparing two entities through a common attribute). It is available in two settings: a few-shot distractor setting where models are provided with 10 paragraphs including the gold ones, and an open-domain full-wiki setting where models must retrieve relevant passages from the entire Wikipedia corpus given only the question.
  • 2WikiMultihopQA (2wiki): A multi-hop question answering dataset that contains complex questions requiring reasoning over multiple Wikipedia paragraphs. Each question is designed to necessitate logical connections across different pieces of information to arrive at the correct answer.
  • Musique (musique): A challenging multi-hop QA dataset containing approximately 25K 2–4 hop questions, constructed by composing single-hop questions from five existing single-hop QA datasets. It is designed to feature diverse and complex reasoning paths, requiring models to integrate information from multiple hops to generate correct answers. The dataset emphasizes comprehensive evaluation of multi-step reasoning capabilities.
@@ C.2 Broad Reasoning Datasets
The following datasets are curated from the UltraDomain (qian2025memorag) benchmark. The benchmark construction leverages financial reports, legal contracts, and 428 college textbooks across 18 distinct domains to evaluate model versatility and adaptability in specialized and broad application scenarios:
@@ Appendix D Baselines
  • NaiveRAG (gao2023rag_survey): A standard chunk-based retrieval baseline that segments raw texts into chunks and stores them in a vector database using text embeddings. For queries, it generates vectorized representations to directly retrieve text chunks based on semantic similarity.
  • GraphRAG (edge2024local): A graph-enhanced RAG system that utilizes an LLM to extract entities and relationships from text, representing them as nodes and edges. It generates community summaries through graph clustering and employs both local (entity-based) and global (community-based) retrieval strategies for comprehensive information access.
  • LightRAG (guo2024lightrag): A graph-structured RAG framework that employs a dual-level retrieval system combining low-level entity retrieval with high-level knowledge discovery. It integrates graph structures with vector representations for efficient retrieval of related entities and their relationships.
  • ToG-2 (ma2024tog2): A knowledge graph-based framework implements a tight-coupling hybrid RAG paradigm that iteratively retrieves information from both unstructured texts and structured knowledge sources. It alternates between graph retrieval and context retrieval for in-depth knowledge exploration.
  • HippoRAG-2 (gutiérrez2025HippoRAG2): A non-parametric continual learning framework that leverages Personalized PageRank algorithm over an open knowledge graph constructed using LLM-extracted triples. It enhances multi-hop reasoning capabilities through sophisticated graph traversal and passage integration mechanisms.
@@ Appendix E Metrics
For Deep Reasoning Tasks, we follow standard QA evaluation practices as ToG (sun2023tog; ma2024tog2) and HippoRAG (hipporag; gutiérrez2025HippoRAG2):
  • Exact Match (EM): Measures the percentage of predictions that exactly match the ground truth answer. Specifically, we follows the Substring-based EM metric (used in ToG/ToG-2 (sun2023tog; ma2024tog2)) to robustly assess answer accuracy in longer, natural-language response generated by LLMs, which goes through the whole response to check whether the answer is in.
@@
For Broad Reasoning Tasks, we adopt a multi-dimensional LLM-based evaluation approach due to the complexity and open-ended nature of these queries following LightRAG (guo2024lightrag):
@@
The LLM-based evaluation uses GPT-4o-mini as judge, with careful attention to prompt design and answer ordering to avoid positional bias. The LLM evaluation prompt is shown in Appendix.L
@@ F.2 Result Detail in Braod Reasoning Tasks
Comprehensiveness 26.1% 73.9% 30.1% 69.9% 10.1% 89.9% 32.5% 67.5%
Diversity 16.9% 83.1% 29.7% 70.3% 7.3% 92.7% 25.9% 74.1%
Empowerment 27.2% 72.8% 30.5% 69.5% 10.1% 89.9% 36.2% 63.8%
Overall 26.3% 73.7% 31.0% 69.0% 9.0% 91.0% 33.7% 66.3%
@@
Comprehensiveness 44.5% 55.5% 47.3% 52.7% 47.3% 52.7% 49.3% 50.7%
Diversity 42.1% 57.9% 46.1% 53.9% 44.5% 55.5% 49.7% 50.3%
Empowerment 22.9% 77.1% 40.9% 59.1% 27.3% 72.7% 36.1% 63.9%
Overall 45.3% 54.7% 46.9% 53.1% 46.1% 53.9% 48.9% 51.1%
@@
Comprehensiveness 36.6% 63.4% 43.3% 56.7% 31.3% 68.7% 45.3% 54.7%
Diversity 29.7% 70.3% 39.7% 60.3% 25.7% 74.3% 37.0% 63.0%
Empowerment 38.2% 61.8% 43.7% 56.3% 31.3% 68.7% 49.7% 50.3%
Overall 37.3% 62.7% 43.7% 56.3% 30.1% 69.9% 47.3% 52.7%
@@
Comprehensiveness 22.2% 77.8% 29.3% 70.7% 19.3% 80.7% 27.3% 72.7%
Diversity 16.5% 83.5% 25.7% 74.3% 15.0% 85.0% 21.4% 78.6%
Empowerment 25.5% 74.5% 30.6% 69.4% 19.3% 80.7% 31.7% 68.3%
Overall 23.3% 76.7% 29.7% 70.3% 18.1% 81.9% 29.3% 70.7%